# A complete classification of the zero-dimensional homogeneous spaces under determinacy

### v2

@@ Abstract.
Chapter 1 Introduction
Throughout this article, unless we specify otherwise, we will be working in the theory $\mathsf{ZF}+\mathsf{DC}$, that is, the usual axioms of Zermelo-Fraenkel (without the Axiom of Choice) plus the principle of Dependent Choices (see Section 2.2 for more details). By space we will mean separable metrizable topological space, unless we specify otherwise. In this chapter, we will give an overview of our results and some of the history that led up to them.
@@ 1.1. History
We begin with the notion that motivates the research contained in this monograph. Intuitively, a space is homogeneous if all of its points “look alike.”
@@
The systematic study of these classes, founded by William Wadge in his doctoral thesis [Wa1] (see also [Wa2]), is known as Wadge theory, and it has become a classical topic in descriptive set theory (see [Ke, Section 21.E]). Under the Axiom of Determinacy, the collection of all Wadge classes on a zero-dimensional Polish space $Z$, ordered by inclusion, constitutes a well-behaved hierarchy that is similar to, but much finer than the well-known Borel hierarchy (and not limited to sets of low complexity). Furthermore, as we will discuss in Section 9.1, when $\mathbf{\Gamma}$ is a good Wadge class in $\omega^{\omega}$ (see Definition 7.8), it will make sense to talk about spaces of exact complexity $\mathbf{\Gamma}$. This concept will feature prominently in the statements (and proofs) of our main results.
In his remarkable doctoral thesis [vE3] (see also [vE1] and [vE2]), Fons van Engelen pioneered the application of Wadge theory to topology by giving a complete classification of the zero-dimensional homogeneous Borel spaces. This classification can be split into three cases. The trivial case is that of the locally compact spaces, since the only zero-dimensional homogeneous ones are the non-empty discrete spaces, $2^{\omega}$, and $\omega\times 2^{\omega}$ (see Proposition 2.14). The next case is that of the low complexity spaces (more precisely, of complexity below $\Delta(\mathsf{D}_{\omega}(\mathbf{\Sigma}^{0}_{2}))$) that are not locally compact. The simplest examples of this kind are $\mathbb{Q}$ and ×$\mathbb{Q}\times 2^{\omega}$. This case can be viewed as a unified generalization of a series of classical results that characterize certain “famous” spaces (these results are neatly collected in [vE1, Theorem 2.4]). The final case is that of the high complexity spaces (more precisely, of complexity above $\Delta(\mathsf{D}_{\omega}(\mathbf{\Sigma}^{0}_{2}))$). It is only in this case that Wadge theory becomes relevant. In fact, the crucial consequence of the high complexity assumption is that it ultimately enabled van Engelen (as it will enable us) to apply a theorem of Steel, and the proof of this implication follows a deeply Wadge-theoretic path.
@@ 1.2. Our results
The aim of this article is to extend van Engelen’s classification beyond the Borel realm, under suitable determinacy assumptions. Naturally, our focus will be on the high complexity case, as the low complexity cases have already been exhaustively treated elsewhere (see [vE1] or [vE3, Chapters 2 and 3]). In particular, under the Axiom of Determinacy, our results will yield a classification of all zero-dimensional homogeneous spaces. Not surprisingly, this classification can be split into a uniqueness result and an existence result, outlined as follows (see Theorem 9.18 for a more precise formulation), where we assume that all spaces are zero-dimensional and $\mathbf{\Gamma}$ is a good Wadge class in $\omega^{\omega}$. This assumption does not result in any loss of generality, since every zero-dimensional homogeneous space of high enough complexity has exact complexity $\mathbf{\Gamma}$ for some good Wadge class $\mathbf{\Gamma}$ in $\omega^{\omega}$ (see Theorem 9.7). We also remark that every homogeneous space is either meager or Baire (see Proposition 2.21). The uniqueness theorem was essentially proved in [CMM1] (see Section 9.3 for more details).
We hope that the reader will not be too irritated by the fact that we have not yet defined most of the relevant notions, as it would not be feasible to do that here. Even so, we believe that stating our main results early will be useful in guiding the reader through this article. We will only say that $\ell(\mathbf{\Gamma})$ and $t(\mathbf{\Gamma})$ respectively denote the level and type of $\mathbf{\Gamma}$. We also remark that at the end of this monograph we give a list of symbols and terminology, which indicates the sections in which they are defined.
@@
$\mathbf{\Gamma}$ has the $\mathbf{\Gamma}$ does not have the separation property separation property $\ell(\mathbf{\Gamma})\geq 2$ There exist both Baire and meager or homogeneous spaces $\ell(\mathbf{\Gamma})=1$ and $t(\mathbf{\Gamma})=3$ of exact complexity $\mathbf{\Gamma}$ There exist only Baire There exist only meager $\ell(\mathbf{\Gamma})=1$ and $t(\mathbf{\Gamma})\in\{1,2\}$ homogeneous spaces homogeneous spaces of exact complexity $\mathbf{\Gamma}$ of exact complexity $\mathbf{\Gamma}$
It will naturally follow from our proof of the above classification result that every zero-dimensional homogeneous space (with the trivial exception of locally compact spaces) is homeomorphic to a semifilter on $\omega$ (see Theorem 9.19). Furthermore, building on the above classification, we will be able to determine precisely for which good Wadge classes $\mathbf{\Gamma}$ in $\omega^{\omega}$ there exists a filter on $\omega$ of exact complexity $\mathbf{\Gamma}$ (see Theorem 10.4). This will allow us to obtain a purely topological characterization of the filters on $\omega$ (see Theorem 10.5). The Borel versions of all the results concerning filters were obtained by van Engelen in [vE5].
Along the way, we will obtain several Wadge-theoretic results of independent interest, culminating with the closure properties of Section 7.5. Even in the case of “folklore” or otherwise known results, we believe that the statements given here will improve on the existing literature in two ways. First, we will generally state our results for uncountable zero-dimensional Polish spaces instead of just $\omega^{\omega}$. Second, we will give level-by-level statements with respect to the determinacy assumptions (see Section 2.2). As noteworthy examples, consider Theorem 5.8 and the results of Section 6.3.
@@
The exposition is self-contained, except for the proofs of fundamental Wadge-theoretic results (for which we refer to [CMM2]), for results pertaining to the low complexity case (for which we refer to [vE3], [vE5] and [Me4]), and for classical results (for which we refer to [En], [Je] and [Ke]).
@@ 1.3. Organization of the paper
In Chapter 2, we will establish our general terminology and notation, and collect several basic results concerning determinacy, the Baire property, filters, semifilters, and homogeneity properties. In Chapter 3, we will give a concise introduction to Wadge theory. As we mentioned above, we will heavily reference [CMM2], although we have tried to make our presentation accessible to readers who are not familiar with this article.
In Chapter 4, we will discuss Hausdorff operations and universal sets, mostly as a tool to construct new Wadge classes (see Corollary 4.16). The intended application is to separated unions (see 6.6). As a bonus, we will show how Hausdorff operations can clarify the notions of level and expansion introduced in Chapter 3.
In Chapter 5, we will discuss the separation property and show that exactly one Wadge class in every non-selfdual pair possesses this property (see Theorem 5.8). In Chapter 6, we will extend Louveau’s notion of type beyond the Borel context, introduce separated unions, then use them to characterize the classes of small type (see Theorem 6.11). These results will be applied in Chapter 7, in order to obtain certain closure properties. For our purposes, the most important results of this kind are those regarding good Wadge classes (see Corollary 7.13) and those that link the level, the type and the separation property with closure under separated unions (see Theorems 7.19 and 7.20).
In Chapter 8, we will give a complete proof of a theorem due to Steel (see Theorem 8.9) which will ultimately provide the homeomorphisms needed to obtain our classification results. An interesting feature of this chapter is that is does not require any determinacy assumptions.
In Chapter 9, we will employ the above results to finally obtain the classification of the zero-dimensional homogeneous spaces. In particular, the closure properties of good Wadge classes will allow us to apply Steel’s Theorem 8.9, in order to obtain the uniqueness part of the classification (see Theorem 9.9). On the other hand, the results on closure under separated unions will allow us to construct the desired homogeneous spaces (see Corollaries 9.14 and 9.15). This chapter begins with a theory of complexity for zero-dimensional spaces, which will enable us to give natural-sounding statements of our classification results (in fact, we have already done so above). This seems to be another worthwhile improvement with respect to van Engelen’s presentation.
In Chapter 10, we will obtain a classification of the filters on $\omega$ up to homeomorphism (see Theorem 10.4). For this purpose, we will employ the notion of closure under squares and prove a further existence result (see Theorem 10.3). Finally, we will discuss the relation between filters and topological groups, including a conjecture that would yield a complete classification of the zero-dimensional topological groups up to homeomorphism (see Conjecture 10.8).
We conclude this section with a diagram that highlights the most significant logical dependencies among the main topics of this article.
${\mathrm{\begin{subarray}{c}\mathrm{Hausdorff}\\ \mathrm{operations}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{\begin{subarray}{c}\mathrm{Level}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$ood ${\mathrm{\begin{subarray}{c}\mathrm{Good}\text{ }\mathrm{Wadge}\\ \mathrm{classes}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{\begin{subarray}{c}\mathrm{Expansions}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{\begin{subarray}{c}\mathrm{Type}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{\begin{subarray}{c}\mathrm{Separated}\\ \mathrm{unions}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{\begin{subarray}{c}\mathrm{Separation}\\ \mathrm{property}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$haracterizationof small ${\mathrm{\begin{subarray}{c}\mathrm{Characterization}\\ \mathrm{of\text{ }small\text{ }type}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{\begin{subarray}{c}\mathrm{Separated}\\ \mathrm{differences}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$reservation of theseparation ${\mathrm{\begin{subarray}{c}\mathrm{Preservation\text{ }of\text{ }the}\\ \mathrm{separation\text{ }property}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$losure propertiesof good Wadge ${\mathrm{\begin{subarray}{c}\mathrm{Closure\text{ }properties}\\ \mathrm{of\text{ }good\text{ }Wadge\text{ }classes}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$reservationof small ${\mathrm{\begin{subarray}{c}\mathrm{Preservation}\\ \mathrm{of\text{ }small\text{ }type}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$losure underseparated ${\mathrm{\begin{subarray}{c}\mathrm{Closure\text{ }under}\\ \mathrm{separated\text{ }unions}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$easonably closedWadge ${\mathrm{\begin{subarray}{c}\mathrm{Reasonably\text{ }closed}\\ \mathrm{Wadge\text{ }classes}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{\begin{subarray}{c}\mathrm{Existence}\\ \mathrm{theorems}\end{subarray}}}$${\mathrm{\begin{subarray}{c}\mathrm{Uniqueness}\\ \mathrm{theorem}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{\begin{subarray}{c}\mathrm{Exact}\\ \mathrm{complexity}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{\begin{subarray}{c}\mathrm{Harrington\text{'}s}\\ \mathrm{lemma}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{\begin{subarray}{c}\mathrm{Steel\text{'}s}\\ \mathrm{theorem}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{\begin{subarray}{c}\mathrm{Knaster\text{-}Reichbach}\\ \mathrm{systems}\end{subarray}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{\begin{subarray}{c}\mathrm{Knaster\text{-}Reichbach}\\ \mathrm{covers}\end{subarray}}}$
Chapter 2 Preliminaries
@@
In this chapter, we will establish much of the terminology and notation to be employed in the remainder of this monograph (with the exception of Wadge theory). We will give basic facts concerning determinacy, the Baire property, filters, semifilters and homogeneity. Furthermore, we will discuss how these notions interact with each other (see for example the diagram in Section 2.4).
@@ 2.1. Terminology and notation
Our reference for basic set theory will be [Je]. Let $\Omega$ be a set. We will denote by $\mathcal{P}(\Omega)$ the collection consisting of all subsets of $\Omega$. When $\Omega$ is a countably infinite set, we will identify every element of $\mathcal{P}(\Omega)$ with its characteristic function, so that every subset of $\mathcal{P}(\Omega)$ will inherit the subspace topology from $2^{\Omega}$. We will denote by $\mathsf{Fin}(\Omega)$ the collection of all finite subsets of $\Omega$, and by $\mathsf{Cof}(\Omega)$ the collection of all cofinite subsets of $\Omega$.
Given $s\in\Omega^{m}$ and $t\in\Omega^{n}$, where $m<\omega$ and $n\leq\omega$, we will denote by $s^{\frown}t\in\Omega^{m+n}$ the concatenation of $s$ and $t$, that is the function defined by setting
$(s^{\frown}t)(i)=\left\{\begin{array}[]{ll}s(i)&\text{if }i<m,\\ t(i-m)&\text{if }m\leq i<m+n,\end{array}\right.$
for $i<m+n$. We will often write $i^{\frown}t$ or $s^{\frown}i$ instead of $(i)^{\frown}t$ or $s^{\frown}(i)$ respectively. We will use the notation $\Omega^{<\omega}=\bigcup_{n<\omega}\Omega^{n}$ and $\Omega^{\leq\omega}=\bigcup_{n\leq\omega}\Omega^{n}$. Given $s\in\Omega^{<\omega}$, we will use the notation $\mathsf{N}_{s}=\{z\in\Omega^{\omega}:s\subseteq z\}$.
Next, assume that $\sigma:\Omega^{<\omega}\longrightarrow\Omega^{<\omega}$ is given. Recall that $\sigma$ is monotone if $\sigma(s)\subseteq\sigma(t)$ whenever $s,t\in\Omega^{<\omega}$ and $s\subseteq t$. We will say that $\sigma$ is eventually growing if $|\sigma(z\upharpoonright n)|\to\infty$ for every $z\in\Omega^{\omega}$. When $\sigma$ is monotone and eventually growing, it makes sense to define $f_{\sigma}:\Omega^{\omega}\longrightarrow\Omega^{\omega}$ by setting
$f_{\sigma}(z)=\bigcup\{\sigma(z\upharpoonright n):n\in\omega\}$
for $z\in\Omega^{\omega}$. For a proof of the following simple fact, see [Ke, Proposition 2.6].
Proposition 2.1.
Let $\Omega$ be a discrete space, and let $f:\Omega^{\omega}\longrightarrow\Omega^{\omega}$. Then the following conditions are equivalent:
  • $f$ is continuous,
  • $f=f_{\sigma}$ for some monotone and eventually growing $\sigma:\Omega^{<\omega}\longrightarrow\Omega^{<\omega}$.
@@
Let $f:X\longrightarrow Y$. Given $S\subseteq X$, we will use the notation $f[S]=\{f(x):x\in S\}$. Given $S\subseteq Y$, we will use the notation $f^{-1}[S]=\{x\in X:f(x)\in S\}$. We will use the notation $\mathsf{dom}(f)=X$ for the domain of $f$, and $\mathsf{ran}(f)=f[X]$ for the range of $f$. When $X=Y$ and $f(x)=x$ for every $x\in X$, we will use the notation $\mathsf{id}_{X}=f$ for the identity on $X$. Recall that $f$ is a retraction if $f$ is continuous, $Y\subseteq X$ and $f\upharpoonright Y=\mathsf{id}_{Y}$.
@@
Our reference for topology will be [En]. A subset of a space is clopen if it is closed and open. A base for a space $X$ is a collection $\mathcal{B}$ consisting of non-empty open subsets of $X$ such that for every $x\in X$ and every non-empty open neighborhood $U$ of $x$ there exists $V\in\mathcal{B}$ such that $x\in V\subseteq U$. A space is zero-dimensional if it is non-empty and it has a base consisting of clopen sets.111 The empty space has dimension $-1$ (see [En, Section 7.1]). A subset $S$ of a space $X$ is codense if $X\setminus S$ is dense. A subset $S$ of a space $X$ is meager if there exist closed nowhere dense subsets $C_{n}$ of $X$ for $n\in\omega$ such that $S\subseteq\bigcup_{n\in\omega}C_{n}$. A subset $S$ of a space $X$ is comeager if $X\setminus S$ is meager. A space is meager if it is a meager subset of itself. A space $X$ is Baire if every non-empty open subset of $X$ is non-meager. Let $X$ and $Y$ be spaces. We will write $X\approx Y$ to mean that $X$ and $Y$ are homeomorphic. An embedding is a function $j:X\longrightarrow Y$ such that $j:X\longrightarrow j[X]$ is a homeomorphism. We will say that an embedding is dense if it has dense range. A function $f:X\longrightarrow Y$ is closed if $f[C]$ is a closed subset of $Y$ whenever $C$ is a closed subset of $X$. Similarly define an open function. The straightforward proof of the following result is left to the reader.
@@
Let $X$ be a metric space, and let $\mathsf{d}$ denote the metric on $X$. Given $x\in X$ and $\varepsilon>0$, we will use the notation $\mathsf{B}(x,\varepsilon)=\{z\in X:\mathsf{d}(z,x)<\varepsilon\}$ for the open ball of radius $\varepsilon$ around $x$. Given $S\subseteq X$, we will use the notation $\mathsf{diam}(S)=\mathsf{sup}\{\mathsf{d}(x,y):x,y\in S\}$ for the diameter of $S$. Given $\mathcal{S}\subseteq\mathcal{P}(X)$, we will use the notation $\mathsf{mesh}(\mathcal{S})=\mathsf{sup}\{\mathsf{diam}(S):S\in\mathcal{S}\}$ for the mesh of $\mathcal{S}$. A metric on a space is compatible if it induces the topology of the space.
@@
The classes defined below constitute the so-called difference hierarchy (or small Borel sets). For a detailed treatment, see [Ke, Section 22.E] or [vE3, Chapter 3]. Here, we will only mention that the classes $\mathsf{D}_{\eta}(\mathbf{\Sigma}^{0}_{\xi}(Z))$ are among the simplest concrete examples of Wadge classes (see Section 3.1).
@@
$\mathsf{D}_{\eta}(A_{\mu}:\mu<\eta)=\left\{\begin{array}[]{ll}\bigcup\{A_{\mu}\setminus\bigcup_{\zeta<\mu}A_{\zeta}:\mu<\eta\text{ and }\mu\text{ is odd}\}&\text{if }\eta\text{ is even,}\\ \bigcup\{A_{\mu}\setminus\bigcup_{\zeta<\mu}A_{\zeta}:\mu<\eta\text{ and }\mu\text{ is even}\}&\text{if }\eta\text{ is odd.}\end{array}\right.$
@@
We conclude this section by stating a classical result that will serve as a fundamental technical tool in Chapter 7 (see [Ke, Theorem 22.16]).222 The conclusion of this theorem could be stated more succinctly by saying that $\mathbf{\Sigma}^{0}_{\xi}(X)$ has the generalized reduction property (see [Ke, Definition 22.14]).

### v3

@@ Abstract.
Throughout this article, unless we specify otherwise, we will be working in the theory $\mathsf{ZF}+\mathsf{DC}$, that is, the usual axioms of Zermelo-Fraenkel (without the Axiom of Choice) plus the principle of Dependent Choices (see §2.2 for more details). By space we will always mean separable metrizable topological space.
1. Introduction
@@ 1.1. History
We begin with the notion that motivates the research contained in this article. Intuitively, a space is homogeneous if all of its points “look alike.”
@@
The systematic study of these classes, founded by William Wadge in his doctoral thesis [Wa1] (see also [Wa2]), is known as Wadge theory, and it has become a classical topic in descriptive set theory (see [Ke, §21.E]). Under the Axiom of Determinacy, the collection of all Wadge classes on a zero-dimensional Polish space $Z$, ordered by inclusion, constitutes a well-behaved hierarchy that is similar to, but much finer than the well-known Borel hierarchy (and not limited to sets of low complexity). Furthermore, as we will discuss in §6.1, when $\mathbf{\Gamma}$ is a good Wadge class in $\omega^{\omega}$ (see Definition 5.8), it will make sense to talk about spaces of exact complexity $\mathbf{\Gamma}$. This concept will feature prominently in the statements (and proofs) of our main results.
In his remarkable doctoral thesis [vE3] (see also [vE1] and [vE2]), Fons van Engelen pioneered the application of Wadge theory to topology by giving a complete classification of the zero-dimensional homogeneous Borel spaces. This classification can be split into three cases. The trivial case is that of the locally compact spaces, since the only zero-dimensional homogeneous ones are the non-empty discrete spaces, $2^{\omega}$, and $\omega\times 2^{\omega}$ (see Proposition 2.13). The next case is that of the low complexity spaces (more precisely, of complexity below $\Delta(\mathsf{D}_{\omega}(\mathbf{\Sigma}^{0}_{2}))$) that are not locally compact. The simplest examples of this kind are $\mathbb{Q}$ and $\mathbb{Q}\times 2^{\omega}$. This case can be viewed as a unified generalization of a series of classical results that characterize certain “famous” spaces (these results are neatly collected in [vE1, Theorem 2.4]). The final case is that of the high complexity spaces (more precisely, of complexity above $\Delta(\mathsf{D}_{\omega}(\mathbf{\Sigma}^{0}_{2}))$). It is only in this case that Wadge theory becomes relevant, ultimately because of the closure properties that it provides. For example, in order to construct a homogeneous space of a given complexity $\mathbf{\Gamma}$, we will start with an arbitrary set of complexity $\mathbf{\Gamma}$, then perform certain operations on it to ensure homogeneity: suitable closure properties will guarantee that these operations preserve complexity (see the proof of Theorem 6.12).
@@ 1.2. Our results
The aim of this article is to extend van Engelen’s classification beyond the Borel realm, under suitable determinacy assumptions. Naturally, our focus will be on the high complexity case, as the low complexity cases have already been exhaustively treated elsewhere (see [vE1] or [vE3, Chapters 2 and 3]). In particular, under the Axiom of Determinacy, our results will yield a classification of all zero-dimensional homogeneous spaces. Not surprisingly, this classification can be split into a uniqueness result and an existence result, outlined as follows (see Theorem 6.17 for a more precise formulation), where we assume that all spaces are zero-dimensional and $\mathbf{\Gamma}$ is a good Wadge class in $\omega^{\omega}$. This assumption does not result in any loss of generality, since every zero-dimensional homogeneous space of high enough complexity has exact complexity $\mathbf{\Gamma}$ for some good Wadge class $\mathbf{\Gamma}$ in $\omega^{\omega}$ (see Theorem 6.7). We also remark that every homogeneous space is either meager or Baire (see Proposition 2.20). The uniqueness theorem was essentially proved in [CMM1] (see §6.2 for more details).
We hope that the reader will not be too irritated by the fact that we have not yet defined most of the relevant notions, as it would not be feasible to do that here. Even so, we believe that stating our main results early will be useful in guiding the reader through this article. We will only say that $\ell(\mathbf{\Gamma})$ and $t(\mathbf{\Gamma})$ respectively denote the level and type of $\mathbf{\Gamma}$. We also remark that right before the bibliography we give a list of symbols and terminology, which indicates the subsections in which they are defined.
@@
$\mathbf{\Gamma}$ has the $\mathbf{\Gamma}$ does not have the
separation property separation property
$\ell(\mathbf{\Gamma})\geq 2$ There exist both Baire and meager
or homogeneous spaces
$\ell(\mathbf{\Gamma})=1$ and $t(\mathbf{\Gamma})=3$ of exact complexity $\mathbf{\Gamma}$
There exist only Baire There exist only meager
$\ell(\mathbf{\Gamma})=1$ and $t(\mathbf{\Gamma})\in\{1,2\}$ homogeneous spaces homogeneous spaces
of exact complexity $\mathbf{\Gamma}$ of exact complexity $\mathbf{\Gamma}$
It will naturally follow from our proof of the above classification result that every zero-dimensional homogeneous space (with the trivial exception of locally compact spaces) is homeomorphic to a semifilter on $\omega$ (see Theorem 6.18). Furthermore, building on the above classification, we will be able to determine precisely for which good Wadge classes $\mathbf{\Gamma}$ in $\omega^{\omega}$ there exists a filter on $\omega$ of exact complexity $\mathbf{\Gamma}$ (see Theorem 7.4). As a corollary, we will obtain a purely topological characterization of the filters on $\omega$ (see Theorem 7.5). The Borel versions of all the results concerning filters were obtained by van Engelen in [vE5], and our generalizations rely heavily on his ideas.
Along the way, we will obtain several Wadge-theoretic results of independent interest, culminating with the closure properties of §5.5. In this regard, the treatment of type given in §4, the preservation results of §5.3 and §5.4, and the theory of complexity described in §6.1 also seem worth mentioning. We remark that, as in [CMM2], we will generally state our results for uncountable zero-dimensional Polish spaces instead of just $\omega^{\omega}$ (which is usually the only case considered in the literature). Furthermore, we will give level-by-level statements with respect to the determinacy assumptions (see §2.2).
@@
While being comfortable with the content of [CMM2] would undoubtedly be helpful, we tried to make the presentation accessible to readers that are not familiar with this article. More specifically, in §3, we collected all the fundamental Wadge-theoretic results that will be needed later on. In addition to being convenient for the reader, this will allow us to incorporate more general versions of known results, with a view towards our intended applications. Furthermore, giving precise statements (and rigorous proofs) of results such as Corollary 3.24, Corollary 3.39, Theorem 3.43 and Theorem 3.50 seems to be valuable beyond the specific scope of this paper.
@@ 1.3. Organization of the paper
In §2, we will establish much of the terminology and notation to be employed in the remainder of this article (with the exception of Wadge theory). We will give basic facts concerning determinacy, the Baire property, filters, semifilters and homogeneity. Furthermore, we will discuss how these notions interact with each other (see for example the diagram in §2.4). As we explained above, §3 is essentially a concise introduction to Wadge theory, which heavily references [CMM2] for the proofs.
In §4, we will extend Louveau’s notion of type beyond the Borel context, introduce separated unions, then use them to characterize the classes of small type (see Theorem 4.11). These results will be applied in §5, in order to obtain certain closure properties. For our purposes, the most important results of this kind are those regarding good Wadge classes (see Corollary 5.13) and those that link the level, the type and the separation property with closure under separated unions (see Theorems 5.19 and 5.20).
In §6, we will employ the above results to finally obtain the existence (and non-existence) results. In conjunction with the results of [CMM1], this will yield the desired classification of the zero-dimensional homogeneous spaces. In particular, the results on closure under separated unions will allow us to construct the required homogeneous spaces (see Corollaries 6.13 and 6.14). This section begins with a theory of complexity for zero-dimensional spaces, based on the closure properties of good Wadge classes, which will enable us to give natural-sounding statements of our classification results (in fact, we have already done so above). This seems to be another worthwhile improvement with respect to van Engelen’s presentation.
In §7, we will obtain the classification and characterization of the filters on $\omega$ mentioned above. For this purpose, we will employ the notion of closure under squares and prove a further existence result (see Theorem 7.3). Finally, we will discuss the relation between filters and topological groups, including a conjecture that would yield a complete classification of the zero-dimensional topological groups up to homeomorphism (see Conjecture 7.8).
2. Preliminaries
@@ 2.1. Terminology and notation
Our reference for basic set theory will be [Je]. Let $\Omega$ be a set. We will denote by $\mathcal{P}(\Omega)$ the collection consisting of all subsets of $\Omega$. When $\Omega$ is countably infinite, we will identify every element of $\mathcal{P}(\Omega)$ with its characteristic function, so that every subset of $\mathcal{P}(\Omega)$ will inherit the subspace topology from $2^{\Omega}$. We will denote by $\mathsf{Fin}(\Omega)$ the collection of all finite subsets of $\Omega$, and by $\mathsf{Cof}(\Omega)$ the collection of all cofinite subsets of $\Omega$. We will use the notation $\Omega^{<\omega}=\bigcup_{n<\omega}\Omega^{n}$. Recall that $\sigma:\Omega^{<\omega}\longrightarrow\Omega^{<\omega}$ is monotone if $\sigma(s)\subseteq\sigma(t)$ whenever $s,t\in\Omega^{<\omega}$ and $s\subseteq t$. Given $s\in\Omega^{<\omega}$, we will use the notation $\mathsf{N}_{s}=\{z\in\Omega^{\omega}:s\subseteq z\}$.
@@
Let $f:X\longrightarrow Y$. Given $S\subseteq X$, we will use the notation $f[S]=\{f(x):x\in S\}$. Given $S\subseteq Y$, we will use the notation $f^{-1}[S]=\{x\in X:f(x)\in S\}$. When $X=Y$ and $f(x)=x$ for every $x\in X$, we will use the notation $\mathsf{id}_{X}=f$ for the identity on $X$. Recall that $f$ is a retraction if $f$ is continuous, $Y\subseteq X$ and $f\upharpoonright Y=\mathsf{id}_{Y}$.
Our reference for topology will be [En]. A base for a space $X$ is a collection $\mathcal{B}$ consisting of non-empty open subsets of $X$ such that for every $x\in X$ and every non-empty open neighborhood $U$ of $x$ there exists $V\in\mathcal{B}$ such that $x\in V\subseteq U$. A local base for a space $X$ at a point $x\in X$ is a collection $\mathcal{B}$ consisting of open neighborhoods of $x$ such that for every open neighborhood $U$ of $x$ there exists $V\in\mathcal{B}$ such that $V\subseteq U$. A $\pi$-base $\mathcal{B}$ for a space is a collection $\mathcal{B}$ consisting of non-empty open sets such that for every non-empty open set $U$ there exists $V\in\mathcal{B}$ such that $V\subseteq U$.
@@
A subset of a space is clopen if it is closed and open. A space is zero-dimensional if it is non-empty and it has a base consisting of clopen sets.222 The empty space has dimension 1-1 (see [En, §7.1]). When we say that a space $X$ is a topological group, we really mean that there exists a topological group whose underlying space is $X$. A subset $S$ of a space $X$ is meager if there exist closed nowhere dense subsets $C_{n}$ of $X$ for $n\in\omega$ such that $S\subseteq\bigcup_{n\in\omega}C_{n}$. A subset $S$ of a space $X$ is comeager if $X\setminus S$ is meager. A space is meager if it is a meager subset of itself. A space $X$ is Baire if every non-empty open subset of $X$ is non-meager.
Let $X$ and $Y$ be spaces. We will write $X\approx Y$ to mean that $X$ and $Y$ are homeomorphic. An embedding is a function $j:X\longrightarrow Y$ such that $j:X\longrightarrow j[X]$ is a homeomorphism. We will say that an embedding is dense if it has dense range. A function $f:X\longrightarrow Y$ is closed if $f[C]$ is a closed subset of $Y$ whenever $C$ is a closed subset of $X$. Similarly define an open function. The straightforward proof of the following result is left to the reader.
@@
The classes defined below constitute the so-called difference hierarchy (or small Borel sets). For a detailed treatment, see [Ke, §22.E] or [vE3, Chapter 3]. Here, we will only mention that the classes $\mathsf{D}_{\eta}(\mathbf{\Sigma}^{0}_{\xi}(Z))$ are among the simplest concrete examples of Wadge classes (see §3.1).
@@
[…]
@@
We conclude this subsection by stating a classical result that will serve as a fundamental technical tool in §5 (see [Ke, Theorem 22.16]).333 The conclusion of this theorem could be stated more succinctly by saying that $\mathbf{\Sigma}^{0}_{\xi}(X)$ has the generalized reduction property (see [Ke, Definition 22.14]).